Red short-horned cattle are homozygous for the red allele, white cattle are homozygous for the white allele, and roan cattle are heterozygotes. Population A consists of 36% red, 16% white, and 48% roan cattle. What is the allele frequency of the red allele? (enter as decimal- zero point something)
Answer #1

Here, the road cattle is heterozygous.
So Let, 'R' is the dominant allele responsible for the red cattle and 'r' is the corresponding recessive allele responsible for the white cattle.
Then, the following are the possible genotypes and corresponding phenotypes -
RR = red cattle
Rr = roan cattle
rr = white cattle

Now, let 'p' be the frequency of the red allele(R) and 'q' be the frequency of the white allele (r).
Also, 'p2' is the frequency of the homozygous red individuals (RR)
'q2' is the frequency of the homozygous white individuals (rr)
and '2pq' is the frequency of heterozygous individuals(Rr).
Here,
p2 (RR) = 36/100 = 0.36
q2 (rr) = 16/100 = 0.16
2pq (Rr) = 48/100 = 0.48

Now, we have to find the frequency of the red allele (R) i.e we have to find the value of p.
p (R) = 0.36 + 0.48/2 = 0.36 + 0.24 = 0.60

Hence, the correct answer is 0.60 .

Please hit on the like button if you are satisfied with the answer. Give comments for further clarification/assistance. Thanks! Have a good day.